The velocity of the body moving in a straight line with a constant acceleration is 10 m/s at a certain instant ''t'' after 5 s the velocity become 20 m/s , the velocity 3 s before  ''t" was????

Grade:9

1 Answers

Khimraj
5 years ago
Let the constant acceleration is a
then
10 + a*5 = 20
so a = 2m/s2
then velocity 3s before t = 10 -2*3 = 4 m/s
Hope it clears...........................................